OpenStudy (anonymous):

solve 16- 2(8w - 7) = 3w - 18(2 - w)

OpenStudy (across):

\[16w-2(8w-7)=3w-18(2-w)\]\[16w-16w+14=3w-36+18w\]\[50=21w\]\[w=\frac{50}{21}\]